Determine if the differential equation (sin y − y sin x)dx + (cos x + x cos y − y)dy = 0 is exact. If it is exact, solve it.

1
Expert's answer
2021-09-06T15:57:56-0400

"\displaystyle
\text{Let $u = (\sin y - y\sin x)dx$ and $v= (\cos x+x\cos y - y)dy$}\\
\frac{\partial (\sin y - y \sin x)}{\partial y} = \cos y - \sin x\\
\frac{\partial (\cos x - x \cos y-y)}{\partial x} = \cos y - \sin x\\
\text{Next, we integrate u with respect to x}\\
\int (\sin y- y\sin x)dx = x\sin y + y \cos x + h(y)-(1)\\
\text{Next, we differentiate the expression above with respect to y}\\
x\cos y + \cos x + h'(y)\\
\text{Comparing the expression above to v, we have}\\
x\cos y + \cos x + h'(y) = \cos x+x\cos y - y\\
\implies h'(y) = -y \therefore h(y) = -\frac{y^2}{2}\\
\text{Substituting h(y) in (1), we have that}\\
x\sin y + y \cos x-\frac{y^2}{2}\\
\text{Hence the the solution of the differential equation is}\\
x\sin y + y \cos x-\frac{y^2}{2}=c"